# OneVision-Encoder: Codec‑Aligned Sparsity as a Foundational Principle for Multimodal Intelligence

### v2

@@ 4.3 Attentive Probing Evaluation
Model Setup Video Benchmarks (Acc. %) Method Backbone Res. Avg. SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 8 Frames / 2048 Patches CLIP ViT-L/14 224 50.5 48.2 46.6 52.2 10.8 52.8 36.1 79.3 78.0 SigLIP ViT-L/16 256 50.1 50.7 43.9 48.9 10.9 52.2 39.1 78.2 77.0 MetaCLIP ViT-L/14 224 48.5 50.6 28.9 49.8 10.4 54.1 37.1 79.6 77.1 MetaCLIP2 ViT-L/14 224 50.2 47.2 48.0 47.7 11.0 48.0 40.9 82.4 76.3 AIMv2 ViT-L/14 224 53.8 55.1 43.6 55.1 12.0 56.6 45.6 81.1 81.3 SigLIP2 ViT-L/16 256 53.1 52.6 50.1 52.7 11.6 54.2 43.8 80.9 79.1 DINOv3 ViT-L/14 224 58.0 57.4 58.6 59.3 13.2 62.5 51.7 82.9 78.6 OV-Encoder (Frame) ViT-L/14 224 58.4 57.7 57.6 58.3 12.1 61.4 52.5 84.3 83.1 OV-Encoder (Codec) ViT-L/14 224 60.2 58.5 67.2 60.0 12.3 62.3 53.9 84.4 83.4 16 Frames / 4096 Patches SigLIP ViT-L/16 256 52.8 52.7 54.7 51.0 11.7 54.1 40.2 79.1 78.8 MetaCLIP2 ViT-L/14 224 51.0 49.3 42.1 51.1 11.2 49.2 43.2 84.0 78.2 AIMv2 ViT-L/14 224 56.4 57.2 55.7 56.4 12.4 58.3 46.2 82.2 82.6 SigLIP2 ViT-L/16 256 55.7 58.2 56.7 53.3 11.9 56.4 45.2 82.7 81.2 DINOv3 ViT-L/14 224 59.1 58.3 61.3 60.8 14.0 63.2 51.9 83.9 79.7 OV-Encoder (Frame) ViT-L/14 224 59.9 58.7 63.2 60.3 12.6 62.9 54.5 85.1 81.6 OV-Encoder (Codec) ViT-L/14 224 61.4 60.1 69.4 60.9 12.9 63.3 54.4 85.4 85.3
@@
Setting Video Benchmarks (Acc. %) AVG SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 Patch Budget = 512 SigLIP2 47.0 42.8 37.3 22.5 46.4 42.9 37.4 74.5 72.4 OV-Encoder (Codec) 55.1 50.0 46.5 50.5 50.6 50.3 41.7 76.7 74.7 Patch Budget = 1024 SigLIP2 53.4 49.0 48.5 35.2 50.0 50.6 42.0 78.8 76.6 OV-Encoder (Codec) 62.4 56.6 54.9 61.6 57.4 58.2 48.4 81.8 80.3 Patch Budget = 2048 SigLIP2 56.6 53.1 52.6 50.1 52.7 54.2 43.8 80.9 79.1 OV-Encoder (Codec) 66.5 60.2 58.5 67.2 60.0 62.3 53.9 84.4 83.4 Patch Budget = 4096 SigLIP2 59.9 55.7 58.2 56.7 53.3 56.4 45.2 82.7 81.2 OV-Encoder (Codec) 67.5 61.4 60.1 69.4 60.9 63.3 54.4 85.4 85.3
@@ 4.4 Patch-Efficient Video Understanding Comparison
Setting Video Benchmarks (Acc. %) AVG SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 OV-Encoder (Codec) 61.4 60.1 69.4 60.9 12.9 63.3 54.4 85.4 85.3 Non-motion Patch Replacement (50%) (Same Video, Same Position) 55.4 52.1 55.4 54.9 11.6 56.3 50.2 83.1 79.4 Counterfactual Motion Replacement (50%) (Motion Patch from Other Video) 54.9 50.6 57.2 53.8 11.3 55.1 49.6 82.7 79.0 Patch–Position Shuffle (Content Preserved, Positions Shuffled) 48.1 41.8 46.3 45.1 8.7 48.2 42.5 78.4 73.6

### v3

@@ 4.3 Attentive Probing Evaluation
Model Setup Video Benchmarks (Acc. %) Method Backbone Res. Avg. SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 8 Frames / 2048 Patches CLIP ViT-L/14 224 50.5 48.2 46.6 52.2 10.8 52.8 36.1 79.3 78.0 SigLIP ViT-L/16 256 50.1 50.7 43.9 48.9 10.9 52.2 39.1 78.2 77.0 MetaCLIP ViT-L/14 224 48.5 50.6 28.9 49.8 10.4 54.1 37.1 79.6 77.1 MetaCLIP2 ViT-L/14 224 50.2 47.2 48.0 47.7 11.0 48.0 40.9 82.4 76.3 AIMv2 ViT-L/14 224 53.8 55.1 43.6 55.1 12.0 56.6 45.6 81.1 81.3 SigLIP2 ViT-L/16 256 53.1 52.6 50.1 52.7 11.6 54.2 43.8 80.9 79.1 DINOv3 ViT-L/14 224 58.0 57.4 58.6 59.3 13.2 62.5 51.7 82.9 78.6 OV-Encoder (Frame) ViT-L/14 224 58.4 57.7 57.6 58.3 12.1 61.4 52.5 84.3 83.1 OV-Encoder (Codec) ViT-L/14 224 60.2 58.5 67.2 60.0 12.3 62.3 53.9 84.4 83.4 16 Frames / 4096 Patches SigLIP ViT-L/16 256 52.8 52.7 54.7 51.0 11.7 54.1 40.2 79.1 78.8 MetaCLIP2 ViT-L/14 224 51.0 49.3 42.1 51.1 11.2 49.2 43.2 84.0 78.2 AIMv2 ViT-L/14 224 56.4 57.2 55.7 56.4 12.4 58.3 46.2 82.2 82.6 SigLIP2 ViT-L/16 256 55.7 58.2 56.7 53.3 11.9 56.4 45.2 82.7 81.2 DINOv3 ViT-L/14 224 59.1 58.3 61.3 60.8 14.0 63.2 51.9 83.9 79.7 OV-Encoder (Frame) ViT-L/14 224 59.9 58.7 63.2 60.3 12.6 62.9 54.5 85.1 81.6 OV-Encoder (Codec) ViT-L/14 224 61.5 60.1 69.4 60.9 12.9 63.3 54.4 85.4 85.3
@@
Setting Video Benchmarks (Acc. %) AVG SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 Patch Budget = 512 SigLIP2 43.4 42.8 28.1 38.7 10.1 42.9 37.4 74.5 72.4 OV-Encoder (Codec) 50.1 50.0 46.5 50.5 10.6 50.3 41.7 76.7 74.7 Patch Budget = 1024 SigLIP2 50.8 49.0 48.7 50.1 10.8 50.6 42.0 78.8 76.6 OV-Encoder (Codec) 56.2 56.6 54.9 58.6 11.1 58.2 48.4 81.8 80.3 Patch Budget = 2048 SigLIP2 53.1 52.6 50.1 52.7 11.6 54.2 43.8 80.9 79.1 OV-Encoder (Codec) 60.2 58.5 67.2 60.0 12.3 62.3 53.9 84.4 83.4 Patch Budget = 4096 SigLIP2 55.7 58.2 56.7 53.3 11.9 56.4 45.2 82.7 81.2 OV-Encoder (Codec) 61.5 60.1 69.4 60.9 12.9 63.3 54.4 85.4 85.3
@@ 4.4 Patch-Efficient Video Understanding Comparison
Setting Video Benchmarks (Acc. %) AVG SSV2 Diving48 Perception Test CharadesEgo Epic-Verb Epic-Noun Kinetics-400 HMDB51 OV-Encoder (Codec) 61.5 60.1 69.4 60.9 12.9 63.3 54.4 85.4 85.3 Non-motion Patch Replacement (50%) (Same Video, Same Position) 55.4 52.1 55.4 54.9 11.6 56.3 50.2 83.1 79.4 Counterfactual Motion Replacement (50%) (Motion Patch from Other Video) 54.9 50.6 57.2 53.8 11.3 55.1 49.6 82.7 79.0 Patch–Position Shuffle (Content Preserved, Positions Shuffled) 48.1 41.8 46.3 45.1 8.7 48.2 42.5 78.4 73.6